# Feedback-Aware Monte Carlo Tree Search for Efficient Information Seeking in Goal-Oriented Conversations

### v1

@@ 3.4 Monte Carlo Tree Search
where $R_{a}(v)$ is the accumulated reward, $N_{v}$ is the visit count of node $v$, $N_{p}$ is its parent node’s visit count, and $C$ is the exploration constant.
@@ 3.5.1 Cluster Assignment and Creation
When a new sample arrives, its text embedding is created and compared with existing cluster medoids to determine its assignment. A similarity threshold $\tau$ is used to decide whether to assign the embedding to an existing cluster or create a new one. Let $\mathbf{e}$ represent the embedding of the current data point and $\mathbf{m}_{k}$ denote the medoid of cluster $k$. The similarity condition is given by:
@@
where $\cdot$ represents the dot product, and $\|\cdot\|_{2}$ represents the L2 norm. If no existing cluster satisfies this condition, a new cluster is created with $\mathbf{e}$ as its initial medoid. If assigned to an existing cluster, $\mathbf{e}$ is added to that cluster’s embeddings, and the medoid $\mathbf{m}_{k}$ is recomputed as:
@@ 3.5.2 Bonus Rewards for Feedback Propagation
For each node $v$, we maintain a dictionary $B_{k}(v)$ which maps each cluster $k$ to the bonus reward (initialized as zero) earned on reaching the target successfully. Once a target is correctly identified, $B_{k}(v)$ is updated for each node along the path from the current node back to the root. $B_{k}(v)$ is a proportion of the node’s cumulative reward and depends on the absolute depth of the node, given by $d_{v}$ controlled by an exponential decay as we move upward in the tree. The bonus reward for node $v$ corresponding to cluster $k$ is updated after every success:
@@
where $\beta$ is a task-specific scaling factor (denoting proportion of total reward), and $\gamma\in(0,1)$ is a decay factor that controls the influence of the bonus, keeping it higher for early-stage questions which have proven effective to narrow down the large possibility set, and lower for questions near the terminal nodes which are often more specific to each case.
To incorporate these bonus rewards into decision-making, we modify the UCT formula used in the selection step to handle similar cases while maintaining exploration capabilities. We introduce a cluster-specific bonus term to bias selection toward nodes associated with successful outcomes for similar data points. The modified UCT formula becomes:
@@
This modification prioritizes nodes from clusters associated with successful outcomes when processing samples similar to their medoids. Algorithm 1 shows the end-to-end summary of the proposed framework.
@@
5:     $\mathcal{C}_{k}\leftarrow\text{AssignCluster}(\mathbf{e}_{s},\mathcal{C},\tau)$triangleright Use Eq. 11-12
@@
9:         if $t<\delta\cdot T$ and $|\Omega_{v}\geq 2|$ then
10:              $v^{*}\leftarrow\text{MCTS}(v,K,\mathcal{C}_{k})$triangleright Use Eq. 4-9, 14
@@
Model Method $\Omega$-aware DX MedDG FloDial
@@
MISQ 87.50 1.97 0.03 89.20 3.46 0.60 74.50 10.15 4.10
MISQ-HF 99.03 2.19 0.05 90.30 3.42 0.41 72.54 10.36 2.94
Table 1: Results on Medical Diagnosis and Troubleshooting Domain in a Closed Set scenario. $\Omega$-aware denotes whether the Questioner Model was informed about the possibility space $\Omega$ once at the beginning of the conversation.
4 Experiments
In this section, we provide an overview of our experiments, including descriptions of the datasets, baselines, evaluation metrics, and implementation details.
4.1 Datasets
We aim to demonstrate the effectiveness and versatility of our approach by focusing on three diverse domains: Medical Diagnosis, Troubleshooting, and 20 Questions. We use the following datasets pre-processed by Hu et al. (2024).
In Medical Diagnosis, a patient initially reports a brief description of their symptoms, based on which the doctor asks questions to diagnose the disease. The maximum number of turns $T$ was limited to 6 in the experiments. Two datasets were used. The DX dataset by Xu et al. (2019) contains 104 doctor-patient dialogues and five diseases in its test set. The MedDG dataset, which originally included over 17,000 conversations across 15 disease types, was refined by removing inconsistent samples. We used 454 high-quality samples for evaluation. Open-ended responses are allowed in MedDG to test the system’s generalization capabilities in less constrained scenarios. Both datasets limit interactions to 6 turns. In the Troubleshooting domain, customer support technicians interact with users to identify faults in systems such as cars or electronic devices. We use the FloDial dataset by Raghu et al. (2021), containing 153 dialogues across 153 unique fault types. Maximum number of turns $T$ was limited to 20. In the 20 Questions domain, the task involves identifying a target item by asking up to 20 yes-or-no questions. The Common dataset includes 111 items spanning categories such as animals, places, food, and objects, and the Things dataset Hebart et al. (2019), was filtered to include 200 items.
4.2 Baselines
The proposed system is compared with two primary baselines: Direct Prompting (DP) and the Uncertainty of Thoughts (UoT) framework Hu et al. (2024). DP involves querying the LLM directly to generate the next question without leveraging any structured planning mechanism. This approach evaluates the basic capability of the LLM to ask effective questions without additional search strategies. It serves as a minimal-effort benchmark, highlighting the improvements achieved by strategic decision-making approaches. UoT implements a tree-based planning method that performs an exhaustive expansion of all possible question paths to maximize information gain. While UoT demonstrates strong performance in reducing uncertainty and improving task success rates, its reliance on full tree exploration up to a predefined depth can lead to significant computational overhead, particularly within large possibility spaces. This baseline is included to measure the improvement in efficiency achieved by the proposed MISQ-HF approach, which selectively expands promising branches rather than exploring all possibilities exhaustively. Comparisons with methods such as Chain-of-Thought (CoT) Wei et al. (2022) and Tree-of-Thoughts (ToT) Yao et al. (2023) are omitted, as prior work by Hu et al. (2024) has demonstrated that UoT consistently outperforms these techniques in both efficiency and effectiveness.
4.3 Evaluation
We use three key metrics to assess both the effectiveness and efficiency of the question-asking system. First, we measure Success Rate (SR), defined as the percentage of cases where the system correctly identifies the target item within the maximum allowed turns. Second, we track Mean Conversation Length in Successful Cases (MSC), which represents the average number of turns needed to reach the correct answer (target) in successful interactions. Finally, we introduce a novel metric to track Question Generation Calls (QGC), defined as the number of LLM prompting calls required for question generation. Lower values indicate better computational efficiency. This metric directly measures the system’s ability to minimize expensive LLM operations while maintaining performance. Together, these metrics provide a comprehensive view of the system’s performance, balancing success rate, conversation efficiency, and computational overhead.
4.4 Experimental Setup
Models. We employed three different LLMs of varying sizes as the Questioner. Llama 3.3 70B Instruct Grattafiori et al. (2024) and Mixtral 8*7B Instruct Jiang et al. (2024) were accessed via the AWS Bedrock API (2023). GPT-4o was accessed via API provided by OpenAI et al. (2024). The user (Answerer) was simulated by Llama 3.3 70B Instruct in all tasks. We prompted the model with ground truth details (e.g., the patient’s disease or fault description) and maintained a separate conversation history. The temperature was set to 0.
MCTS implementation. We set the number of iterations $K=10$ and exploration constant $C=0.2$. The maximum simulation depth $d_{s}$ was set to 3 to balance computational efficiency with search effectiveness. For each $\Omega_{v}$, the question generator creates $m=3$ potential questions to maintain diversity. For the reward calculation in $R_{IG}(v)$, the scaling parameter $\lambda$ was set to 0.4, for both MCTS and UoT.
Feedback mechanism. We used a decay factor $\gamma=0.9$ for the bonus rewards, ensuring early-stage questions receive higher weights. The cluster similarity threshold $\tau$ was set to 0.9 in terms of cosine similarity, and the bonus scaling factor $\beta$ was set to 0.2 for all tasks. Problem descriptions were embedded using DistilBERT Sanh et al. (2019) for the troubleshooting domain, and Clinical-BERT Wang et al. (2023) for medical diagnosis. To optimize computational efficiency, we cache the decision tree, allowing reuse across all samples within each dataset. The system tracks Mean QGC through a thread-safe counter.
@@ 5 Results and Discussion
Table 1 shows our results in the Closed Set scenario. Our approach requires fewer QGC while maintaining higher SR compared to baseline methods, indicating both improved effectiveness and computational efficiency. Notably, the $\Omega$-aware Questioner consistently performs better as compared to when unaware. To demonstrate the importance of the feedback mechanism, we also report results using MISQ, our framework without the hierarchical feedback component, to rationalize the design choice and support an ablation study.
In the Medical Diagnosis domain, MISQ-HF achieves superior performance across both datasets, with an average reduction of 6.77 times in Mean QGC across all three LLMs. Using Llama 3.3 70B as a Questioner, MISQ-HF required only 0.04 Mean QGC compared to UoT’s 0.77 on DX, showing a 19.25 times reduction. Similarly, a reduction of 8.97 times is observed on MedDG. This efficiency stems from selective tree expansion rather than exhaustive exploration. In the medical domain, the cost of misdiagnosis far outweighs the overhead of a few additional clarifying questions. While MSC is marginally higher for MISQ-HF compared to UoT, this trade-off can be overseen by the higher success rate of diagnosis, which directly impacts patient outcomes.
In the Troubleshooting domain using the FloDial dataset, MISQ-HF continues to excel with higher SR, surpassing UoT by 20.68% improvement on average across the three LLMs. Our framework also achieves notable computational efficiency as compared to UoT, with an average of 18.63 times reduction in Mean QGC. While troubleshooting conversations typically require more turns due to fault complexity, MISQ-HF maintains a similar MSC (a difference of less than 3 turns on average) while delivering improved performance. With GPT-4o, MISQ-HF achieved an SR of 72.54% with 27 clusters, the largest consisting of 31 samples.
Table 2 shows performance on the 20 Questions domain, evaluated using the Common and Things datasets. Due to a lack of problem description or initial hints about the target, we do not evaluate MISQ-HF on this domain. For both datasets, MISQ consistently outperforms UoT and DP in all three metrics. These results demonstrate the scalability of the proposed approach to larger possibility spaces.
@@
Table 2: Results on 20 Questions Datasets in the Closed Set scenario

### v2

@@ 3.4 Monte Carlo Tree Search (MCTS)
where $R_{total}(v)$ refers to the cumulative reward, $N_{v}$ is the visit count of node $v$, $N_{p}$ is its parent node’s visit count, and $C$ is the exploration constant. Note that $R_{total}(v)$ is initialized to 0 in the beginning and updated by adding expected reward $R_{e}(v)$ over time.
@@ 3.5.1 Cluster Assignment and Creation
When a new sample (user) arrives, we generate a text embedding of their problem description. This embedding is then compared with existing cluster medoids to determine its assignment. A similarity threshold $\tau$ is used to decide whether to assign the embedding to an existing cluster or create a new one. Let $\mathbf{e}$ represent the embedding of the current data point and $\mathbf{m}_{k}$ denote the medoid of cluster $k$. The similarity condition is given by:
@@
where $\cdot$ represents the dot product, and $\|\cdot\|_{2}$ represents the L2 norm. If no existing cluster satisfies this condition, a new cluster is created with $\mathbf{e}$ as its initial medoid. If assigned to an existing cluster, $\mathbf{e}$ is added to that cluster’s set of embeddings $\mathcal{C}_{k}$, and the medoid $\mathbf{m}_{k}$ is recomputed as:
@@ 3.5.2 Bonus Rewards for Feedback Propagation
For each node $v$, we maintain a dictionary $B_{k}(v)$ which maps each cluster $k$ to the bonus reward (initialized as zero) earned on reaching the target successfully. Once the target ($\omega$) is correctly identified, $B_{k}(v)$ is updated for each node along the path from the current node back to the root, if $\omega\in\Omega_{v}$. $B_{k}(v)$ is a proportion of the node’s cumulative reward and depends on the absolute depth of the node, given by $d_{v}$ controlled by an exponential decay as we move upward in the tree. The bonus reward for node $v$ corresponding to cluster $k$ is updated after every success:
@@
where $\beta$ is a task-specific scaling factor (denoting proportion of total reward), and $\gamma\in(0,1)$ is a decay factor that controls the influence of the bonus. $B_{k}(v)$ is a scalar value that is higher for early-stage questions which have proven effective to narrow down the large possibility set, and lower for later-stage questions near the terminal nodes which are often more specific to each unique case.
To incorporate these bonus rewards into decision-making, we modify the UCT formula used in the selection step to handle similar cases while maintaining exploration capabilities. We introduce a cluster-specific bonus term to bias selection toward question nodes that led to successful outcomes for similar data points in the past. The modified UCT formula becomes:
@@
4 Experiments
In this section, we provide an overview of the datasets, baselines, evaluation metrics, and implementation details. Algorithm 1 shows the pseudocode of the proposed framework.
4.1 Datasets
We demonstrate the effectiveness of our approach by focusing on three diverse domains: Medical Diagnosis (MD), Troubleshooting (TS), and 20 Questions. We use the following datasets pre-processed by [7]. In Medical Diagnosis, a patient initially reports a brief description of their symptoms, based on which the doctor asks questions to diagnose the disease. The maximum number of turns $T$ was limited to 6 in the experiments. Two datasets were used. The DX dataset [19] contains 104 doctor-patient dialogues and five diseases in its test set. The MedDG dataset, which originally included over 17,000 conversations across 15 disease types, was refined by removing inconsistent samples. We used 454 high-quality samples for evaluation. Open-ended responses are allowed in MedDG to test the system’s generalization capabilities in less constrained scenarios. Both datasets limit interactions to 6 turns. In the Troubleshooting domain, customer support technicians interact with users to identify faults in systems such as cars or electronic devices. We use the FloDial dataset [13], containing 153 dialogues across 153 unique fault types. Maximum number of turns $T$ was limited to 20. In the 20 Questions domain, the task involves identifying a target item by asking up to 20 yes-or-no questions. The Common dataset includes 111 items spanning categories such as animals, places, food, and objects, and the Things dataset [6], was filtered to include 200 items.
@@ 4.1 Datasets
5:    $\mathcal{C}_{k}\leftarrow\text{AssignCluster}(\mathbf{e}_{s},\mathcal{C},\tau)$triangleright Use Eq. 10-11
@@
9:         if $t<\delta*T$ and $|\Omega_{v}\geq 2|$ then
10:             $v^{*}\leftarrow\text{MCTS}(v,K,\mathcal{C}_{k})$triangleright Use Eq. 3-8, 13
@@
4.2 Baselines
We evaluate two primary baselines: Direct Prompting (DP) and the Uncertainty of Thoughts (UoT) framework [7]. DP directly queries the LLM to generate the next question without structured planning, serving as a minimal-effort benchmark to highlight improvements achieved by strategic decision-making. UoT uses tree-based planning to expand question paths that maximize information gain. While effective in reducing uncertainty, its exhaustive tree exploration can be expensive. This baseline assesses the efficiency improvement of the proposed MISQ-HF approach, which selectively expands promising branches. Comparisons with Chain-of-Thought (CoT) [18] and Tree-of-Thoughts (ToT) [20] are omitted, as prior work [7] shows UoT consistently outperforms these methods.
4.3 Evaluation
We use three key metrics to evaluate our system’s effectiveness and efficiency. First, Success Rate (SR) measures the percentage of cases where the system correctly identifies the target within the maximum allowed turns. Second, Mean Conversational length in Successful Cases (MSC) to track the number of turns required. Finally, we introduce a novel metric to track Question Generation Calls (QGC), defined as the number of LLM prompting calls required for question generation during planning.
4.4 Experimental Setup
Models. We employed three different LLMs of varying sizes as the Questioner. Llama 3.3 70B Instruct [4] and Mixtral 8*7B Instruct [9] were accessed via the AWS Bedrock [1]. GPT-4o was accessed via API from OpenAI [11]. The user (Answerer) was simulated by Llama 3.3 70B Instruct in all tasks. We prompted the model with ground truth details (e.g., the patient’s disease or fault description) and maintained a separate conversation history. The temperature was set to 0. The ratio of turns was defined by $\delta=0.6$.
MCTS implementation. We set the number of iterations $K=10$ and exploration constant $C=0.2$. Maximum simulation depth $d_{s}$ was set to 3 to balance computational efficiency with search effectiveness. For each $\Omega_{v}$, the LLM was prompted to generate $m=3$ potential questions to maintain diversity. For the reward calculation in $R_{IG}(v)$, the scaling parameter $\lambda$ was set to 0.4. Experiments were run on an 8-core CPU with 16 GB RAM.
Feedback mechanism. We used a decay factor $\gamma=0.9$ for the bonus rewards. The cluster similarity threshold $\tau$ was set to 0.9 in terms of cosine similarity, and the bonus scaling factor $\beta$ was set to 0.2 for all tasks. Problem descriptions were embedded using DistilBERT [14] for the troubleshooting domain, and Clinical-BERT [17] for medical diagnosis. To optimize computational efficiency, we cache the decision tree, allowing reuse across all samples within each dataset. The system tracks Mean QGC through a thread-safe counter.
Table 1: Results on MD and TS domains in a Closed Set scenario. $\Omega$-aware denotes whether the Questioner Model was informed about the possibility space $\Omega$ at the beginning of the conversation.
Model Method $\Omega$-aware MD: DX MD: MedDG TS: FloDial
@@ 4.4 Experimental Setup
MISQ 87.50 1.97 0.05 89.20 3.46 0.60 74.50 10.15 4.10
MISQ-HF 99.03 2.19 0.03 90.30 3.42 0.41 72.54 10.36 2.94
Table 2: Results on MD and TS Domain in a Closed Set scenario, when the root node is initialized with the constrained set of possibilities $\Omega_{c}\subseteq\Omega$. Leads to improvement in SR and MSC.
Model $\Omega$-aware MD: DX MD: MedDG TS: FloDial
[…]
Llama 3.3 70B Instruct $\times$ 85.57 0.93 0.20 80.61 2.81 6.13 43.13 6.83 28.26
98.07 0.96 0.21 89.20 2.67 4.09 72.54 8.25 23.35
Mixtral 8*7B Instruct 90.38 1.48 0.24 89.87 2.27 4.69 59.47 7.18 30.30
GPT-4o 99.03 0.54 0.05 92.95 2.15 3.69 78.43 7.18 13.47
@@ 5 Results and Discussion
Table 1 shows results in the Closed Set scenario. Our approach requires fewer QGC while maintaining higher SR compared to baseline methods, indicating both improved effectiveness and computational efficiency. Notably, the $\Omega$-aware Questioner consistently performs better as compared to when unaware. To demonstrate the importance of the feedback mechanism, we also report the results using MISQ, our framework without the hierarchical feedback component, to rationalize the design choice and support an ablation study.
Figure 2: Gain in SR when using MISQ-HF initialized with a constrained set.
Across both datasets in the MD domain, MISQ-HF achieves superior performance, with an overall average reduction of 6.77 times in Mean QGC over UoT. Notably, using Llama 3.3 70B, MISQ-HF required only 0.04 Mean QGC compared to UoT’s 0.77 on DX, showing a 19.25 times reduction. Similarly, a reduction of 8.97 times is observed on MedDG. This efficiency stems from selective tree expansion rather than exhaustive exploration. In the medical domain, the cost of misdiagnosis far outweighs the overhead of a few additional clarifying questions. While MSC is marginally higher for MISQ-HF compared to UoT, this trade-off can be overseen by the higher success rate of diagnosis, which directly impacts patient outcomes.
In the TS domain using FloDial dataset, MISQ-HF continues to excel with higher SR, surpassing UoT by 20.68% improvement on average across the three LLMs. Our framework also achieves notable computational efficiency as compared to UoT, with an average of 18.63 times reduction in Mean QGC. While troubleshooting conversations typically require more turns due to fault complexity, MISQ-HF maintains a similar MSC (a difference of less than 3 turns on average) while delivering improved performance.
Table 2 reports the performance when we initialized the root as a constrained set of possibilities. This generates more relevant questions early on, based on the user’s problem description. Figure 2 shows the notable gain in SR using GPT-4o. On average, across the three LLMs on MD datasets, we find that SR is improved by 27.52% over DP, 11.91% over UoT and 5.58% over MISQ-HF initialized with full superset. Across both domains, an average of $\approx 8\%$ gains were observed using constrained set. This comes with a trade-off in QGC, as the tree gets wider and increases the planning requirements. Additional analysis, example conversations and prompts are provided in the Appendix.
Table 3: Results on 20 Questions Data: Closed Set
@@
Table 3 shows performance on the general information-seeking domain, evaluated using 20-Questions on Common and Things datasets. Due to a lack of problem description or initial hints about the target, we cannot evaluate MISQ-HF here. For both datasets, MISQ consistently outperforms UoT and DP in all three metrics. These results demonstrate the scalability of the proposed approach to larger possibility spaces.
Results in the Open Set scenario are provided in Table 5 of Appendix. Our extensive evaluation shows that the MISQ-HF outperforms other baselines on datasets with different sizes of possibility spaces. The hierarchical feedback mechanism enables the system to learn from successful questioning trajectories, especially in domains like troubleshooting and medical diagnosis, where similar cases often recur. Performance without hierarchical feedback (MISQ) further emphasizes the benefits of this approach. Compared to exhaustive methods like UoT (Table 1), MISQ-HF achieves $\approx$10x reduction in Mean QGC while delivering $\approx 12\%$ improvement in SR across these domains.
Limitations. The system currently does not incorporate a mechanism to learn from mistakes in failure cases, which could potentially refine its decision-making process and adapt to edge cases more effectively. Penalizing redundant or suboptimal quality of questions also requires careful design of reward function and remains an open challenge. Incorporating confidence metrics to better quantify uncertainty in risk-sensitive domains is another potential improvement. Addressing these limitations presents a promising avenue for future research.
6 Conclusion
In this paper, we introduced Monte Carlo Tree Search for Information Seeking Questions with Hierarchical Feedback (MISQ-HF), a novel framework that addresses fundamental limitations in adaptive question-asking for goal-oriented conversational systems. Our approach makes three key contributions to the literature on conversational planning: (1) a principled inference-time planning procedure for selecting optimal questions via selective tree expansion, (2) a hierarchical feedback mechanism that incorporates historical performance signals in the UCT formulation to modulate exploration-exploitation trade-offs, and (3) an effective and efficient method that reduces the number of LLM calls made during the planning phase. Empirical validation across domains such as medical diagnosis, technical troubleshooting, and general information seeking demonstrates substantial improvements in task performance, while reducing and resource utilization by an order of magnitude. The observed gains in constrained possibility spaces further suggest promising avenues for integration with domain-specific knowledge structures.
The broader implications of our work extend beyond immediate performance metrics to address fundamental challenges in human-machine collaborative problem-solving. By combining structured planning with adaptation mechanisms derived from historical interactions, we establish a foundation for conversational AI systems capable of reasoning under uncertainty. Future directions include extending MISQ-HF to multi-agent settings, integrating explainability into planning, and investigating theoretical convergence under varying domain characteristics. We posit that MISQ-HF represents a significant step toward conversational systems that can engage in truly adaptive, efficient, and goal-directed information acquisition—a critical capability for next-generation AI assistants.
Appendix
Appendix A Improvement in Success Rate with $\Omega_{c}$
Figure 3 demonstrates the % Change reported in Table 4. We find that an additional 8% gain in success rate is observed on average when we start with a constrained set of possibilities as compared to full superset of possibilities using MISQ-HF. Even larger gains are observed as compared to baselines UoT and DP.
Figure 3: Improvement in Success Rate on MD and TS Domain in a Closed Set scenario, when initializing the root node with the constrained set of possibilities $\Omega_{c}\subseteq\Omega$.
Table 4: Improvement in Success Rate on MD and TS Domain in a Closed Set scenario, when using MISQ-HF-c, that is, initialization of root node with the constrained set of possibilities $\Omega_{c}\subseteq\Omega$. MISQ-HF-s denotes initialization from full superset at the root node.
Model Dataset DP UoT MISQ-HF-s MISQ-HF-c
% Change
over DP
% Change
over UoT
% Change over
MISQ-HF-s
GPT-4o DX 73.07 82.69 99.03 99.03 35.52 19.76 0.00
MedDG 81.27 88.79 90.3 92.95 14.37 4.68 2.93
FloDial 43.79 59.47 72.54 78.43 79.10 31.88 8.12
Llama 3.3 70B Instruct DX 88.46 79.8 98.07 98.07 10.86 22.89 0.00
MedDG 84.14 89.86 93.39 89.2 6.01 -0.73 -4.48
FloDial 21.56 60.78 67.97 72.54 236.45 19.35 6.72
Mixtral 8*7B Instruct DX 50.0 76.92 70.19 90.38 80.76 17.49 28.77
MedDG 76.43 83.7 84.58 89.87 17.58 7.371 6.25
FloDial 16.99 39.21 49.01 59.47 250.03 51.67 21.34
Appendix B Open Set Results